Sunday 03/18/12

11:00 A – 8:45 P

  • No colonies on plates
  • Finish O/N STORM run.  Run failed at movie 121, unknown error message.
  • attempt bead feducial drift correction:
  • polysterene beads good autofluoresnce in all but 750, but bleach fast.
  • 405 beads not visible in other channels past 488.
  • Bcd/Hb and Ftz stained embryos incubating 1 hr in primary at RT
  • image new antibody stains, results recorded in staining record:
  • Summary in brief: en, AbdB, and HP1 work after standard fix and MeOH storage.
  • Running test gel of PCR linearization of Hox genes and en-locus genes to do new in vitro Tx reactions
  • Fly work: flipped all GFP lines into 2 bottles each.
  • Fix embryos 15 min, test DSHB antibodies on shorter freshly fixed embryos:  m a-Psc, m a-Esc, m a-eve,
  • Heat shock transform new cells with old ligation.  Mix cells myself, troubleshoot control for KE error.
  • Heat fix MP09 embryos (try to match protocol better, have ice at ready).
  • Rpt transformations at 37 on partioned plate O/N with IPTG and X-GAL.
  • Need to make more amp plates!
Images from antibody testing: 

This entry was posted in Summaries and tagged , , , , , , . Bookmark the permalink.